OpenStudy (anonymous):

Solve using the square root method. x2 = 64

OpenStudy (anonymous):

What happens when you square root 64?

OpenStudy (anonymous):

8

OpenStudy (maaz):

8

OpenStudy (this):

8*8=64

OpenStudy (anonymous):

Don't forget:\[\sqrt{x^2}=\sqrt{64} \color{blue}\Longrightarrow x = \color{red} \pm 8\]